# Explicitly Modeling Subcortical Vision with a Neuro-Inspired Front-End Improves CNN Robustness

### v1

@@ E.2 SubcorticalBlock Implementation
Center Radius [deg] [Solomon338] 0.042 0.063 0.042 0.066
Surround Radius [deg] [Solomon338] 0.279 0.602 0.162 0.565
Excitation Radius [deg] [Solomon338] 0.236 0.289 0.070 0.116
Inhibition Radius [deg] [Solomon338] 0.564 0.869 0.312 0.411
Suppression Index [Solomon338] 0.808 0.719 0.813 0.539
Saturation Index [RaghavanENEURO.0515-22.2023] 0.095 0.365 0.200 0.470
@@
When defining the search space for each variable in our Bayesian optimization framework, our primary objective was to minimize the introduction of inductive biases by employing search spaces as broad as feasible. In many cases, this was straightforward — for instance, we constrained parameters like the semisaturation constant, $c_{\text{50}}$ to lie within physically meaningful bounds. However, for parameters such as the center and surround radii of the DoG filters, the radius of the Gaussian used in the contrast normalization layer, and the ratio of peak contrast sensitivity, we adopted a more heuristic approach. Specifically, we drew on values reported in the neuroscience literature to inform the bounds of the search space. For the DoG center and surround radii, we defined symmetric search intervals of centered around values reported in the reference study to which we aimed to maximize alignment [Solomon338]. Similarly, the bounds for the normalization radius were guided by reported relationships between the suppressive field and the surround Gaussian [Bonin2005-cu]. The same principle was applied to the contrast sensitivity ratio [CRONER19957]. Table E6 provides a comprehensive overview of all parameters tuned, including the corresponding search bounds and the literature references used to guide their selection, where applicable.
Table E6: Hyperparameter search space for the SubcorticalBlock. Minimum and maximum bounds used in the Bayesian optimization for each hyperparameter of the SubcorticalBlock: center radius of the DoG ($r_{c}$); surround radius of the DoG ($r_{s}$); peak contrast sensitivity ratio ($k_{s}/k_{c}$); light adaptation nonlinearity constant ($\gamma$); contrast normalization pooling radius ($r_{\text{CN}}$); semisaturation constant ($c_{50}$); and contrast normalization exponent ($n$). For a subset of these hyperparameters, literature references were used to inform the choice of search bounds.
@@
Layer Hyperparam. Min. Max. Min. Max. Reference
Light Adapt. $\gamma$ 0.01 2.0 0.01 2.0
DoG Conv. $r_{\text{c}}$ [deg] 0.034 0.050 0.050 0.76 Solomon et al. [Solomon338]
$r_{\text{s}}$ [deg] 0.223 0.335 0.482 0.722 Solomon et al. [Solomon338]
$k_{\text{s}}/k_{\text{c}}$0.068 -0.003 -0.037 -0.002 Croner and Kaplan [CRONER19957]
Contrast $r_{\text{CN}}$ [deg] 0.140 0.419 0.301 0.903 Bonin et al. [Bonin2005-cu]
Normalization $c_{\text{50}}$ 0.01 1.0 0.01 1.0
$n$ 0.01 1.0 0.01 1.0
@@
We employed Bayesian optimization using the gp_minimize function from the Scipy library [2020SciPy]. The optimization was performed over a defined parameter space for 640 evaluations, with 64 initial points generated using a Sobol sequence. The acquisition function was probabilistically selected among Lower Confidence Bound (LCB), Expected Improvement (EI), and Probability of Improvement (PI) at each iteration. The exploration-exploitation balance was controlled using $\kappa=1.96$ for LCB and $\xi=0.01$ for EI and PI.
E.3 Training Details
All models were trained on an internal cluster, using a 48GB NVIDIA A40 GPU with Python 3.11 and PyTorch 2.2+cu117, taking roughly tree days to train.
@@ E.3 Training Details
During training, images were randomly horizontally flipped with a probability of 0.5, then resized and cropped to 224$\times$224 pixels. For models without the SubcorticalBlock, images were normalized by subtracting and dividing by [0.5, 0.5, 0.5]. During evaluation, images were resized to 256 pixels on the shorter side, followed by a center crop to 224$\times$224 pixels, and the same normalization was applied.
@@
Models were trained using a cross-entropy loss between ground-truth labels and predicted logits, with label smoothing [7780677] of 0.1. Optimization was performed using stochastic gradient descent with momentum set to 0.9 and weight decay of $5\times 10^{-4}$. Training was conducted for 50 epochs with a batch size of 256. We employed the 1-Cycle learning rate policy [smith2018], where the learning rate was initialized at 0.004, increased to 0.1 at 30% of the total training steps, and then annealed to $4\times 10^{-7}$ following a cosine schedule. When using PRIME [10.1007/978-3-031-19806-9_36], we fine-tuned a standardly trained model for an additional 50 epochs using the same training protocol, except with a maximum learning rate of 0.01 reached at 10% of the training schedule.

### v2

@@ E.2 SubcorticalBlock Implementation
Light adaptation pooling size.
Because the primate visual system exhibits both global and local forms of light adaptation, we initially modeled luminance adaptation as a spatially local process, implemented via Gaussian filtering analogous to our contrast normalization layer. Interestingly, during Bayesian optimization, the learned filter radius consistently expanded to encompass nearly the entire image, suggesting that global rather than local adaptation better supported LGN response property prediciton. To reduce computational overhead, we therefore adopted a global luminance normalization strategy.
@@
Center Radius [deg] [58] 0.042 0.063 0.041 0.064
Surround Radius [deg] [58] 0.279 0.602 0.289 0.620
Excitation Radius [deg] [58] 0.236 0.289 0.094 0.125
Inhibition Radius [deg] [58] 0.564 0.869 0.226 0.609
Suppression Index [58] 0.808 0.719 0.710 0.610
Saturation Index [51] 0.095 0.365 0.200 0.410
@@
When defining the search space for each variable in our Bayesian optimization framework, our primary objective was to minimize the introduction of inductive biases by employing search spaces as broad as feasible. In many cases, this was straightforward — for instance, we constrained parameters like the semisaturation constant, $c_{\text{50}}$ to lie within physically meaningful bounds. However, for parameters such as the center and surround radii of the DoG filters, the radius of the Gaussian used in the contrast normalization layer, and the ratio of peak contrast sensitivity, we adopted a more heuristic approach. Specifically, we drew on values reported in the neuroscience literature to inform the bounds of the search space. For the DoG center and surround radii, we defined symmetric search intervals of centered around values reported in the reference study to which we aimed to maximize alignment [58]. Similarly, the bounds for the normalization radius were guided by reported relationships between the suppressive field and the surround Gaussian [22]. The same principle was applied to the contrast sensitivity ratio [87]. Table E2 provides a comprehensive overview of all parameters tuned, including the corresponding search bounds, the literature references used to guide their selection, where applicable, and the final hyperparameters.
Table E2: SubcorticalBlock hyperparameters and search space used for tuning. Minimum ($x_{\text{min}}$) and maximum ($x_{\text{max}}$) bounds used in the Bayesian optimization for each hyperparameter of the SubcorticalBlock and hyperparameter optima obtained ($x^{\ast}$). Values describe center radius of the DoG ($r_{c}$); surround radius of the DoG ($r_{s}$); peak contrast sensitivity ratio ($k_{s}/k_{c}$); contrast normalization pooling radius ($r_{\text{CN}}$); semisaturation constant ($c_{50}$); contrast normalization exponent ($n$). While not obtained through optimization, the kernel sizes used ($k_{\text{DoG}}$ and $k_{\text{CN}}$) are also presented. For a subset of these hyperparameters, literature references were used to inform the choice of search bounds.
@@
Layer $x$ $x_{\text{min}}$ $_{\text{max}} xx^{\ast}$ $x_{\text{min}}$ $_{\text{max}} xx^{\ast}$ Ref.
DoG $r_{\text{c}}$ [deg] 0.034 0.050 0.047 0.050 0.76 0.76 [58]
Conv. $r_{\text{s}}$ [deg] 0.223 0.335 0.224 0.482 0.722 0.534 [58]
$k_{\text{s}}/k_{\text{c}}$ -0.068 -0.003 -0.12 -0.037 -0.002 -0.004 [87]
$k_{\text{DoG}}$ 19 33
Contrast $r_{\text{CN}}$ [deg] 0.140 0.419 0.419 0.301 0.903 0.902 [22]
Norm. $c_{\text{50}}$ 0.01 1.0 1.0 0.01 1.0 0.19
$n$ 0.01 1.0 1.0 0.01 1.0 0.81
$k_{\text{CN}}$ 43 69
@@
We employed Bayesian optimization using the gp_minimize function from the Scipy library [88]. The optimization was performed over a defined parameter space for 640 evaluations, with 64 initial points generated using a Sobol sequence. The acquisition function was probabilistically selected among Lower Confidence Bound (LCB), Expected Improvement (EI), and Probability of Improvement (PI) at each iteration. The exploration-exploitation balance was controlled using $\kappa=1.96$ for LCB and $\xi=0.01$ for EI and PI.
E.3 EVNet Variants
For all EVNet variants, we re-estimated the scaling factor applied to the VOneBlock whenever it was included, and adjusted the V1 noise Fano factor such that the accumulated Fano factor was 1. Apart from these modifications, most variants were derived by simply performing the modifications described in previous sections, with the two exceptions detailed below.
Contrast normalization ablation.
Because the light adaptation and contrast normalization layers operate in close synchrony, removing the contrast normalization layer substantially destabilized training. In particular, the absence of contrast normalization caused activations within the SubcorticalBlock to explode, primarily due to excessively high responses from the light adaptation mechanism. This effect was most pronounced when image (or image crops) contained small, bright regions surrounded by dark backgrounds — conditions that produce low mean activations but locally high responses in Equation  2. To mitigate this, we modified the light adaptation layer’s mean computation to ignore pixel values below a threshold of $\epsilon=0.05$, effectively preventing spurious amplification of isolated bright pixels.
LGN–V2 skip connection.
When incorporating the skip connections between the SubcorticalBlock and the VOneBlock bottleneck, we maintained a total channel dimensionality of 64 at the input to the backend model. Of these, 60 channels originated from the bottleneck output, while the remaining 4 channels were adapted activation maps from the SubcorticalBlock. Given that the VOneBlock operates with a stride of 4, we applied a $5\times 5$ max-pooling operation with the same stride to the SubcorticalBlock activations prior to concatenation, ensuring spatial alignment and consistent feature scaling across pathways.
E.4 Training Details
All models were trained on an internal cluster, using 48GB NVIDIA A40 GPUs with Python 3.11, PyTorch 2.2 with CUDA 11.7, taking roughly tree days to train.
@@ E.4 Training Details
During training, images were randomly horizontally flipped with a probability of 0.5, then resized and cropped to $\times224$ pixels. Images were normalized by subtracting and dividing by [0.5, 0.5, 0.5], with the exception of model that included the light adaptation layer of the SubcorticalBlock. During evaluation, images were resized to 256 pixels on the shorter side, followed by a center crop to $\times224$ pixels, and the same normalization was applied.
@@
Models were trained using a cross-entropy loss between ground-truth labels and predicted logits, with label smoothing [89] of 0.1. When using the ResNet50 and CORnet-Z architectures, optimization was performed using stochastic gradient descent with momentum set to 0.9 and weight decay of $5\times 10^{-4}$. For EfficientNet-B0, we used RMSProp with a momentum of 0.9, smoothing constant of 0.9, and a denominator stability term of 1.0. Training was conducted for 50 epochs with a batch size of 256. We employed the 1-Cycle learning rate policy [90], where the learning rate was initialized at 4% of the maximum learning rate, increased up to maximum at 30% of the total training steps, and then annealed to $4\times 10^{-4}\%$ of the maximum following a cosine schedule. For the ResNet50, the maximum learning rate was set to 0.1; for the EfficientNet-B0, it was set to 0.256; and, for CORnet-Z, it was set to 0.05. When using PRIME [34], we fine-tuned a standardly trained model for an additional 50 epochs using the same training protocol, except with a maximum learning rate of 0.01 reached at 10% of the training schedule.